Curriculum Vitae

Personal Ezra Haber Glenn, AICP 8 Hall Street Somerville, Massachusetts 02144 eglenn@public-planning.org ; eglenn@mit.edu 617.253.2024 (w) ; 617.628.9881 (h) ; 617.721.7131 (c)

Education MA, 1996 University of California at Davis Social Anthropology; concentration in Sociology of Planning and Policy-Making. Honors: Wakeham Teaching Fellowship (two years), National Science Foundation finalist (honorable mention).

BA, 1991 University of Pennsylvania Anthropology. Honors: Phi Beta Kappa, magna cum laude, Dean’s List, Benjamin Franklin Scholar.

Professional Experience 2002-present Public Planning, Research, and Implementation, Inc.

Founder and Executive Director

2007-present Massachusetts Institute of Technology, Cambridge, Massachusetts Department of Urban Studies and Planning Lecturer and Special Assistant to the Department Head

2006-2007 City of Lawrence, Massachusetts Community Development Department Director of Community Development

2004-2006 City of Somerville, Massachusetts Mayor’s Office of Strategic Planning & Community Development Director of Planning & Development

2003-2004 City of Somerville, Massachusetts Office of Housing & Community Development Director of Commercial & Economic Development

1999-2003 McGregor & Associates, P.C., Boston, Massachusetts Environmental & Land Use Planner

1996-1999 City of Somerville, Massachusetts Planning Department Senior Planner

1996-1997 Lincoln Institute of Land Policy, Cambridge, Massachusetts Research Assistant

1992-1993 Catholic Charities Refugee & Immigration Services, Boston, Massachusetts VISTA Volunteer Teaching Lecturer, “Lawrence Practicum: Information, Asset-building, and the Immigrant

City,” Massachusetts Institute of Technology, Fall 2011 (graduate level).

Lecturer, “Quantitative Reasoning and Statistical Methods for Planning,” Massachusetts Institute of Technology, Spring 2011 (graduate level core required course).

Organizer/Facilitator, “American Urban Planning,” Massachusetts Institute of Technology, Fall 2010, SPURS/Humphrey Seminar.

Lecturer, “Quantitative Reasoning and Statistical Methods for Planning,” Massachusetts Institute of Technology, Spring 2010 (graduate level core required course).

Instructor, “City of Boston” Commonwealth School, Boston, Fall, 2009 (high school short-course).

Organizer/Facilitator, “American Urban Planning,” Massachusetts Institute of Technology, Fall 2009, SPURS/Humphrey Seminar.

Lecturer, “Quantitative Reasoning and Statistical Methods for Planning,” Massachusetts Institute of Technology, Spring 2009 (graduate level core required course).

Lecturer, “Lawrence Practicum: Information, Asset-building, and the Immigrant City,” Massachusetts Institute of Technology, Spring 2009 (graduate level).

Instructor, “City of Boston,” Commonwealth School, Boston, Spring, 2009 (high school short-course).

Lecturer, “New Theories From and For Practice I,” Massachusetts Institute of Technology, Fall 2008 (graduate level).

Adjunct Professor, “Community Mapping,” University of Massachusetts–Lowell, Department of Regional Economic and Social Development, Spring 2007 (graduate level).

Adjunct Professor, “Community Mapping,” University of Massachusetts–Lowell, Department of Regional Economic and Social Development, Spring 2006 (graduate level).

Adjunct Professor, “Urban Planning, Urban Politics, Urban Power,” Tufts University, Experimental College, Spring 2002 (undergraduate level).

Teaching Assistant, “Introduction to Cultural Anthropology,” University of California at Davis, 1994-1996 (three quarters).

Teaching Assistant, “The Phillipines,” University of California at Davis, 1996 (upper level undergraduate).

Teaching Assistant, “Cultural Ecology,” University of California at Davis, 1995 (upper level undergraduate).

Teaching Assistant, “Political Economy,” University of California at Davis, 1995 (upper level undergraduate).

Presentations “Making Use of Local Census Data.” Mel King Institute for Community Building, Cambridge, MA, March 14, 2012.

“inTeractive Somerville, or: How I Learned to Stop Worrying and Love the Internet.” TEDx Somerville, Somerville, MA, March 4, 2012.

“Neighborhood-Level Census Data for Community Development.” Mel King Institute for Community Building/Local Initiatives Support Coalition (LISC), Boston, MA, October 6, 2011.

“Data for Planning: Using the American Community Survey.” Central Massachusetts Regional Planning Agency, Worcester, MA, September 20, 2011.

“acs.R: An R Package for Urban Planners Using Neighborhood-Level Data From the U.S. Census.” Conference for Computers in Urban Planning and Urban Management, Banff, Alberta, July 7, 2011.

“Making Sense of the American Community Survey.” Massachusetts Association of Planning Directors Conference, New Bedford, MA, June 10, 2011.

“Planning for People: How to Make the Community Central to the Process.” Fenway Community Development Corporation Annual Meeting (panel discussion), Boston, MA, May 3, 2011.

“Where is Research Informing Sustainability Planning.” American Planning Association National Conference, Boston, MA, April 9, 2011.

“Dashboards, Mashboards, Washboards, Planning Boards, and Other (Possibly Futile) Attempts to Make Sense of Overwhelming Information in a Complex World.” MIT PlanningTech Conference, Cambridge, MA, April 8, 2011.

“Planning for Community Development Practitioners: Data Analysis, ” Mel King Institute for Community Building training course, Boston, MA, November 19, 2010.

“Quantitative Methods for Planning.” APA/AICP Prep Course, Watertown, MA, May 3, 2010.

“Patterns in Space and Time: Data Visualization for Regional Planning,” Tufts University Urban and Environmental Policy and Planning, Medford, MA, April 14, 2010.

“Planning with Community Support,” Citizen Planner Training Collaborative, Worcester, MA, November 4, 2009.

“The Death and Life of Great American City Planning,” Ignite Boston 6, Boston, MA, September 17, 2009.

“Magic Tricks for Public Meetings,” MIT Independent Activities Period session, Cambridge, MA, January 27, 2009.

“Arts and Economic Development in Somerville, MA,” MIT Independent Activities Period session, Cambridge, MA, January 22, 2009.

“Planning with Community Support,” Citizen Planner Training Collaborative, Fitchburg, MA, November 18, 2008.

“Recommendations for Floodplain Management in R’lyeh,” Symposium on Infrastructure and Disaster Mitigation, Miskatonic University, Arkham, MA, October 31, 2008.

“Planning with Community Support,” Citizen Planner Training Collaborative, Lowell, MA, October 22, 2008.

“Leadership Paradoxes: CDBG and Strategies for Concentrated Neighborhood Revitalization,” U.S. Department of Housing and Community Development Community Development Leadership Institute Training, Boston, MA, May 20, 2008.

“Low-Tech Planning Games,” CoLab Skillz Workshop, Massachusetts Institute of Technology, Cambridge, MA, April 10, 2008.

“Special Permits and Variances,” Citizen Planner Training Collaborative, Danvers, MA, November 15, 2007.

“Planning Games: Beyond SimCity,” American Planning Association Southeastern New England Regional Conference, Hyannis, MA, October 11, 2007.

“Foreclosure Prevention and Predatory Lending” (moderator), National Community Development Association Annual Conference, Burlington, VT, Sept 18, 2007.

“Community/University Partnership: the View from Both Sides of the Fence,” Eighth Annual Committee on Industrial Theory and Assessment Conference, Lowell, MA, April 26, 2007.

“Neighborhood Revitalization Strategies: Making the Most of CDBG,” Citizen Planner Training Collaborative Spring Conference, Worcester, MA, March 17, 2007.

“Affordability and Home Economics in Lawrence, MA: Pre-qualified for Foreclosure,” MassHousing Mortgage Broker training, Framingham, MA, December, 2006.

“EPA Brownfields Mobile Workshop, Lawrence, MA,” EPA Brownfields Conference, Boston, MA, November 14, 2006.

“How to Prepare an Open Space & Recreation Plan,” training presented on three different occasions to the Massachusetts Association of Conservation Commissions, 2002-2004.

“The EOEA Smart Growth Fiscal Impact Tool,” training series at five locations across Massachusetts, 2002-2003.

“APA and Planning History,” American Planning Association Southeastern New England Regional Conference, Providence, RI, 2002.

“Environmental Zoning: Sustaining Resources through Growth Management,” Lincoln Institute of Land Policy, 2000.

“The Transportation-Jobs Connection,” Planner’s Network Annual Conference, Lowell, MA, June 18, 1999.

“411-IRRI: The International Rice Research Institute and the Emerging Information Technologies,” paper presented to session on “Applied Anthropology of Computing” at the Annual Meeting of the Society for Applied Anthropology, March 27-31, 1996, Baltimore.

Presentations before literally hundreds of planning meetings, charettes, visioning sessions, public forums.

Publications Shape Your Neighborhood: Participatory Urban and Regional Planning, O’Reilly Media, forthcoming (2012; tentative title).

Municipal Benchmarks for the Middle Cities: a preliminary look at the data, Pioneer Institute White Paper No. 60, May, 2010.

“Review of Handbook of Research on Urban Informatics: The Practice and Promise of the Real-Time City,” Journal of the American Planning Association, 75(4).

“Review of The Ghost Map: The Story of London’s Most Terrifying Epidemic and How It Changed Science, Cities, and the Modern World,” Journal of the American Planning Association, 74(3).

“The Effects of Performance Measurement on Community Development Spending Priorities: CDBG meets GPRA,” Neighbourhood: the international journal of urban regeneration, 1:1, 2008.

“Planning Students in the Planning Office: the view from both sides of the fence,” paper presented at the Eighth Annual Conference of the Committee on Industrial Theory and Assessment, University of Massachusetts–Lowell, 2007.

“Mapping through History, History through Mapping,” New England Planning, August, 2005.

Consulting Editor, Waste: Refuse, Misuse, and Reuse, “Science@Work” Series, Raintree/Weigl Educational Publishers Ltd., 2004.

Consulting Editor, Ecosystems: Species, Spaces, and Relationships, “Science@Work” Series, Raintree/Weigl Educational Publishers Ltd., 2004.

“Open Space Planners Workbook,” Massachusetts Executive Office of Environmental Affairs, 2003 (co-written with Ralph Willmer, AICP, under contract with the Commonwealth of Massachusetts).

“EOEA Smart Growth Fiscal Impact Tool Manual,” Massachusetts Executive Office of Environmental Affairs, 2003 (co-written with Ralph Willmer, AICP, under contract with the Commonwealth of Massachusetts; also provided theoretical assistance to programming team in the development of associated fiscal impact modeling software tool, and led focus groups and training sessions for community end-users).

Editor, New England Planning (monthly, typically 10-16 pages), 1998-2002.

“Community Preservation Watch,” occasional/semi-regular column in New England Planning, 2000-2002.

“Planning by the Numbers,” occasional/semi-regular column in New England Planning, 1999-2002.

“Guide to the Lower Mystic Basin Land & Water Trail,” Mystic River Watershed Associates, 2000

“Water Resource Planning for Environmental Zoning: Sustaining Resources through Growth Management,” Charles River Watershed Association/Conservation Law Foundation/Lincoln Institute of Land Policy, 1999 (with Pamela DiBona and Bennet Heart)

Research and editorial assistance to Gregor I. McGregor in revising and updating (while employed at McGregor & Associates, P.C., 1998-2003):

- Massachusetts Environmental Law (edited volume, Massachusetts Continuing Legal Education, Inc.)

- Mathew-Bender’s treatise on Brownfields (Massachusetts chapter)

- Mathew-Bender’s treatise on State Environmental Law (Massachusetts chapter)

Editorial and research assistant, Lincoln Institute of Land Policy “Policy Focus Report” series:

- “Open Space Conservation: Investing In Your Community’s Economic Health,” John Tibbetts, 1998.

- “Risks and Rewards of Brownfield Redevelopment,” James Wright, 1997.

Assistant Editor, The Berkeley Guides: The Pacific Northwest and Alaska, Fodor’s, 1994.

Plans and Projects “Affordable Housing Production Plan,” Town of Easton, MA, 2010 (with Community Opportunities Group, Inc.).

“The Arlington Neighborhood: a revitalization strategy,” City of Lawrence, MA, 2009 (with Community Opportunities Group and Land Strategies, Inc.).

“North Common Neighborhood Revitalization Strategy Area Plan,” City of Lawrence, MA, 2006 (with Cahill Associates).

“City of Lawrence Community Development Department: Management/Reorganization Plan,” 2006 (with Cahill Associates).

“East Somerville Neighborhood Revitalization Strategy Area Plan,” Office of Strategic Planning & Community Development, Somerville, MA, 2005.

Open Space & Recreation Plan, Office of Housing & Community Development, Somerville, MA, 2005.

Master Plan, Town of Scituate, MA, 2004 (with McGregor & Associates).

Master Plan, Town of Littleton, MA, 2002 (with Ken Kreutziger, FAICP).

“Groton 2020” (master plan), Town of Groton, MA, 2002.

Master Plan, Town of Weymouth, MA, 2001 (with The Cecil Group, Inc.).

Community Development Plan, Town of Braintree, MA 2001 (with Thomas Planning Services).

Over a dozen master plans, open space plans, and other planning studies and reports (with McGregor & Associates).

Code/Software/Programming acs.R: a package for the R statistical programming language, providing a general toolkit for downloading, managing, and analyzing data from the U.S. Census American Community Survey. (Author and maintainer; current version 0.8:

http://cran.r-project.org/web/packages/acs/index.html.)

EOEA Smart Growth Fiscal Impact Tool: developed theoretical framework for planning tool to estimate fiscal impacts of new development on local budgets; assisted with testing, customization, and training.

Guest Lectures “Introduction to American Planning,” MIT Special Program for Urban and Regional Studies (SPURS), Fall 2011.

“Introduction to American Planning,” MIT Special Program for Urban and Regional Studies (SPURS), Fall 2010.

“Planning Field Projects,” Tufts University Department of Urban and Environmental Planning and Policy, Professor Rusty Russell, Spring 2008 (practicum client).

“MIT@Lawrence Practicum,” Massachusetts Institute of Technology, Department of Urban Studies and Planning, Professors Lorlene Hoyt and Langley Keyes, Fall 2007 (client/guest lecturer).

“MIT@Lawrence Practicum,” Massachusetts Institute of Technology, Department of Urban Studies and Planning, Professors Lorlene Hoyt and Langley Keyes, Fall 2006 (client/guest lecturer).

“Planning Field Projects,” Tufts University Department of Urban and Environmental Planning and Policy, Professor Rusty Russell, Spring 2005 (panelist on planning ethics).

“Brownfields Practicum,” Harvard Graduate School of Design, Department of Landscape Architecture, Professor Niall Kirkwood, Spring 2005 (client/juror).

“Architecture Studio on Union Square, Somerville” Northeastern University Department of Architecture, Professor Christina Crawford, Spring 2005 (client/juror).

“Architecture Studio on Union Square, Somerville” Northeastern University Department of Architecture, Professor Michael LeBlanc, Fall 2004 (client/juror).

“Expanding Boston’s Core: Somerville and Its Competitive Context,” Harvard Graduate School of Design Studio, Professor Francois Vigier, Fall 2004 (client/adviser/juror).

“Central Artery Planning & Design Studio,” Citizen Schools, 2002 (juror).

Thesis Advising “Understanding Unemployment and Local Hiring in Lawrence, Massachusetts: A Report for the City of Lawrence” (MCP thesis advisor), Polina Bakhteiarov, Massachusetts Institute of Technology, 2011.

“The Future of Manufacturing in Somerville” (MCP thesis reader), Anne Emig, Massachusetts Institute of Technology, 2011.

“When the Train Comes: Exploring the Use of Property and Land Acquisition Funds to Ensure Affordability in Future Transit Station Areas” (MCP thesis reader), Sandra Padilla, Massachusetts Institute of Technology, 2010.

“Creating an Asset Management Model for Massachusetts State-Aided Public Housing” (MCP thesis advisor), Hannah H. Creeley, Massachusetts Institute of Technology, 2009. [*Outstanding MCP Thesis Award Winner, Department of Urban Studies & Planning*]

“The Effect of Organizational Capacity on Urban Redevelopment Outcomes: The Case of the Community Redevelopment Agency of Los Angeles and the Baldwin Hills Crenshaw Plaza” (MCP thesis reader), Virginia Flores, Massachusetts Institute of Technology, 2009.

“Land Use and Climate Change in Miami-Dade County” (MCP thesis reader), Haley R. Peckett, Massachusetts Institute of Technology, 2009.

“Evaluating the Massachusetts Historic Rehabilitation Tax Credit as a Tool for Promoting Historic Preservation and Economic Development” (MCP thesis advisor), Kathryn C. Emmitt, Massachusetts Institute of Technology, 2008.

“The Green Line is Coming! Affordable Housing Development and Preservation Strategies for the Somerville Community Corporation” (MCP thesis reader), Cali Kay Gorewitz, Massachusetts Institute of Technology, 2008.

“Out of Bounds?: Rhetoric of Urban Form and Its Influence on State Legislation in Massachusetts” (MCP thesis advisor), Benjamin P. Solomon-Schwartz, Massachusetts Institute of Technology, 2008.

Awards Department of Urban Studies & Planning, Outstanding Contribution to Student Life (2009)

Professional Affiliations and Volunteer Positions Current American Institute of Certified Planners, member

American Planning Association, member; Massachusetts Chapter Board of Directors (1998-2003); MIT University Liaison (2010-present)

Public Planning, Research, and Implementation, Inc., Executive Director and Founder

Somerville Community Corporation, President

Mel King Institute for Community Building: Steering Committee; Curriculum Committee; Academic Roundtable

Executive Order 418 Pre-Approved Planning Consultant, Department of Housing and Community Development, Commonwealth of Massachusettts

Reviewer, Adams Arts Program Implementation Grants, Massachusetts Cultural Council

On-Call Planning Consultant, Massachusetts Housing Partnership 40B Affordable Housing Technical Assistance Program Past Agassiz Cooperative Preschool, Treasurer (2001-2003)

New England Planning, Editor (1998-2002) Mystic River Watershed Association, Board of Directors (1997-2002); Vice President (2000-2002) Alewife/Mystic River Advocates, Board of Directors (1996-2000) Somerville Conservation Commission, Associate Member (1996-1999)

March 19, 2012